# Multi-Exit Kolmogorov–Arnold Networks: enhancing accuracy and parsimony

### v1

@@ Abstract
Kolmogorov–Arnold Networks (KANs) uniquely combine high accuracy with interpretability, making them valuable for scientific modeling. However, it is unclear a priori how deep a network needs to be for any given task, and deeper KANs can be difficult to optimize. Here we introduce multi-exit KANs, where each layer includes its own prediction branch, enabling the network to make accurate predictions at multiple depths simultaneously. This architecture provides deep supervision that improves training while discovering the right level of model complexity for each task. Multi-exit KANs consistently outperform standard, single-exit versions on synthetic functions, dynamical systems, and real-world datasets. Remarkably, the best predictions often come from earlier, simpler exits, revealing that these networks naturally identify smaller, more parsimonious and interpretable models without sacrificing accuracy. To automate this discovery, we develop a differentiable “learning to exit” algorithm that balances contributions from exits during training. Our approach offers scientists a practical way to achieve both high performance and interpretability, addressing a fundamental challenge in machine learning for scientific discovery.
@@ 1 Introduction
Despite these promising qualities, KANs face challenges that can limit their effectiveness in scientific applications. Training KANs presents optimization difficulties, as the learnable univariate functions must be carefully parameterized and refined, with deeper networks often proving particularly challenging to optimize [15]. The architecture search problem—determining the appropriate number of layers and widths—–also remains significant, as practitioners must balance expressiveness against parsimony while avoiding overfitting [20]. Seeking smaller, more parsimonious models without sacrificing accuracy is crucial because overly deep KANs lose interpretability as compositions of many univariate functions become difficult to understand, while smaller KANs better retain the interpretability that makes them valuable for scientific modeling [15]. These challenges suggest that architectural innovations are needed to better realize KANs’ potential for scientific discovery.
@@
Experiments demonstrate the effectiveness of multi-exit KANs across diverse scientific modeling tasks, showing consistent improvements in both accuracy and parsimony compared to traditional single-exit KANs. Our key contributions include: (1) the first application of multi-exit architectures to KANs, with a joint training framework that enables deep supervision; (2) empirical validation across regression problems, dynamical systems modeling, continual learning scenarios, and real-world datasets; (3) evidence that multi-exit KANs often achieve better performance at earlier exits, indicating more parsimonious models without sacrificing accuracy; and (4) a “learning to exit” algorithm that addresses the choice of extra hyperparameters when using multi-exits. Additionally, we provide insights into why multi-exit architectures are particularly well-suited for KANs and discuss the mechanisms underlying their improved performance. These results suggest that multi-exit architectures represent a valuable enhancement to KANs for scientific applications, offering a principled approach to balancing accuracy and interpretability.
The rest of this paper is organized as follows. Section 2 provides background on Kolmogorov–Arnold networks and multi-exit and early-exit neural architectures. Section 3 describes our approach for incorporating multiple exits into the KAN architecture. Section 4 presents results comparing single-exit and multi-exit KANs across multiple domains. Section 5 introduces the learning to exit method. Finally, we conclude in Sec 6 by discussing the implications of our findings, including why multi-exit architectures improve KAN performance, and directions for future work.
@@ 2.1 Kolmogorov–Arnold networks
In a KAN, the activation functions are typically parameterized using B-splines, local polynomial approximations of the one-dimensional functions. Although many other function-fitting techniques have been considered, including radial basis functions [30], Fourier series [31], sinusoidal functions [32], Chebyshev polynomials [33], and wavelet-based representations [34], all of which have many advantages, particularly in terms of computational efficiency, for our purposes here we focus on the original B-spline approach.
@@
To form a deep network by stacking layers of learned activation functions, summation units (and, optionally, multiplication [16]) aggregate the outputs from the previous layer’s activation functions. The number of units across $L$ layers is $[d=N_{0},N_{1},\ldots,N_{L}=m]$. The shape of the KAN is the vector of widths $[N_{0},N_{1},\ldots,N_{L}]$. (See Fig. 1 for an example of a KAN network with shape $[2,3,2,1]$.) Between layers $i$ and $i+1$ there are $N_{i}N_{i+1}$ activation functions. Following [15], $\phi_{\ell,j,i}(x)$ denotes the activation function connecting unit $i$ in layer $\ell$ to unit $j$ in layer $\ell+1$ ($\ell=0,\ldots,L-1$; $i=1,\ldots,N_{\ell}$; $j=1,\ldots,N_{\ell+1}$). The summation units then combine the activation functions to propagate information through the network: for the signal $x_{\ell+1,j}$ into unit $j$ in layer $\ell+1$, giving
@@
KANs are trained with a loss function $\mathcal{L}=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{reg}}$ comprising a data loss and a regularization loss, with the hyperparameter $\lambda$ determining regularization strength. Usually the data loss is mean squared error (MSE):
@@
where $y_{i}$ is the true value for observation $i$ and $\hat{y}_{i}$ is the KAN’s prediction for that observation, while the regularization loss is a combination of an L1 norm and an entropy both defined on the activation functions (for details, see Liu et al. [15]). Activation function parameters are learned via gradient-based optimization to minimize $\mathcal{L}$ on training data. Training commonly includes a refinement process where the resolution of the B-spline grids is increased; gradually increasing grid resolution during training has been shown to improve KAN accuracy better than using a finer grid from the start [15]. KANs are typically optimized with quasi-Newton methods, particularly L-BFGS [35], also employed in this work, though first-order methods such as SGD or Adam [36] can be used as well.
@@ 2.2 Multi-exit and early exit networks
Multi-exit networks are a deep learning architecture where some or all hidden layers in the network have an additional branch point called an exit [21, 22, 23]. These exits are small subnetworks, often only a single layer, whose output can be used alongside or instead of the main trunk network’s output. The most common application is for deep classifiers [21, 22, 37, 38], often for computer vision or language models. In a classification task, the network is typically given inputs of varying difficulties, For easy inputs, i.e., those far from the decision boundary, the network may be able to classify accurately using only basic features built by the earlier layers. But for difficult inputs, the network may need to utilize all the layers to build more complex features in order to make a successful classification [21]. By equipping the network with multiple exits, an easy input can reduce inference-time compute by using the output of an early exit only, saving both time and energy when making predictions [24]. The promise of efficiency gains motivates the study of early exit networks and learning to exit algorithms. This approach goes by various names in the literature, including deeply supervised networks, cascaded learning, conditional deep learning, and adaptive inference. For more on multi-exit and early-exit networks, see Scardapane et al. [24], Laskaridis et al. [39] or Rahmath et al. [40].
@@ 3 Adding exits to KANs
A multi-exit KAN augments a standard (single-exit) KAN of shape $[d=N_{0},N_{1},\ldots,N_{L}=m]$ with additional exits as follows. For each layer $\ell$ of the KAN, beginning from the input and continuing until the second-to-last hidden layer [^1] The last hidden layer already has an exit, the main trunk output., add an exit layer, another KAN network, with shape $[N_{\ell},m]$. The $m$-dimensional output of these exits can then be used to predict the same output as the main trunk exit. We illustrate a multi-exit KAN alongside the corresponding single-exit version in Fig. 1.
@@
where $\mathcal{L}_{k}=\sum_{i}\left(y_{i}-\hat{y}_{i}^{(k)}\right)^{2}/n$ is the MSE for exit $k$ and $w_{k}$ is the weight for exit $k$. The exit weights satisfy $\sum_{k}w_{k}=1$. These weights become a hyperparameter to be tuned by the researcher using validation data and this tuning process in practice has been straightforward. Equation (9) is quite flexible as the exit weights allow us to prioritize certain exits by weighting them more heavily than others, and individual exits can even be disabled by zeroing their weights. However, for a network with many exits, these $K-1$ degrees of freedom become a large search space. Therefore, after our main results, in Sec. 5 we propose and apply a “learning to exit” method to automatically learn $w$ alongside the KAN parameters.
@@
How many more parameters are added to a KAN by adding exits? The number of parameters in a KAN depends on its architecture, which dictates the number of activation functions, and the number of parameters per activation function. The parameters per activation function depends on the number and order of the spline bases (Eq. (5)), assuming B-splines are used to model the activation functions. This is the same for activation functions in the main trunk and in the exits, so we only need to consider the number of activation functions to determine the overhead added to a KAN by adding exits.
The number of activation functions between layers in a standard KAN is the product of the layer widths, so a KAN with shape $[d=N_{0},N_{1},\ldots,N_{L}=m]$ will have
@@
Notice that the sum of layer widths will be smaller than the sum of products of adjacent layer widths, (unless the layers are all one unit), so, unless $m$ is large, the main trunk dominates the number of parameters in the KAN. Indeed, for the case of uniform layer width, $N_{\ell}=d$ for all $\ell$, the main trunk will have $(L-1)d^{2}+dm$ activation functions, including the original exit, and the newly added exits will introduce $(L-1)dm$ activation functions in total. In this case, the new exits will contribute fewer activation functions than the main branch when $m<d$, typical of regression problems (Eq (1)) and a nearly equal number (fewer by $dm$) of activation functions when $m=d$, typical of dynamical systems modeling (Eqs. (2) or (3)).
Also, note that no one prediction made by the model will use all the activation functions, even if they were all used during training. Thus, multi-exits provide their benefits with reasonable, often modest, parameter overhead.
@@ 4 Results
[…]
@@ 4.1 Regression tasks
For the sinc function a KAN of shape $[1,2,2,2,1]$ performed well. See Appendix A for full details on training settings and hyperparameters and data generation. This KAN may seem deep for such a function. KANs can support multiplication units [16], although they are not strictly necessary due to the KART, so we decided to forgo them for simplicity and therefore expected a deeper KAN to perform better for this task, hence the aforementioned shape. Parameterizing the multi-exit weighted loss with a simple linear ramp, $w=[1,2,3,4]$ (unnormalized) worked well: performance on test data was good, with the final exit showing an order of magnitude lower error than the equivalent single-exit network (Fig. 2). In fact, three of the four exits, all but Exit 0, outperformed the single-exit network, indicating robust and parsimonious (parameter-efficient) capture of the data.
@@
[…]
@@
[…]
@@
[…]
@@ 4.2 Data-driven modeling of dynamical systems
where $\phi_{n}=0.4-6\left(1+x_{n}^{2}+y_{n}^{2}\right)^{-1}$ and bifurcation parameter $\mu=0.9$. KANs, unlike sparse regression methods, have been show to model the Ikeda map well [19].
[…]
Figure 4: Multi-step prediction of the Ikeda map (Eq. (14). The multi-exit KAN tracks the dynamics well for about twice as many steps into the future (shaded regions) as the single-exit KAN. (Blue: ground truth; orange: KAN prediction.)
@@
[…]
Figure 5: Multi-step prediction of the ecosystem (Eq (15)). The multi-exit KAN encodes the dynamics well, not diverging significantly until $t>1000$.

### v2

@@ Abstract
Kolmogorov–Arnold Networks (KANs) uniquely combine high accuracy with interpretability, making them valuable for scientific modeling. However, it is unclear a priori how deep a network needs to be for any given task, and deeper KANs can be difficult to optimize and interpret. Here we introduce multi-exit KANs, where each layer includes its own prediction branch, enabling the network to make accurate predictions at multiple depths simultaneously. This architecture provides deep supervision that improves training while discovering the right level of model complexity for each task. Multi-exit KANs consistently outperform standard, single-exit versions on synthetic functions, dynamical systems, and real-world datasets. Remarkably, the best predictions often come from earlier, simpler exits, revealing that these networks naturally identify smaller, more parsimonious and interpretable models without sacrificing accuracy. To automate this discovery, we develop a differentiable “learning-to-exit” algorithm that balances contributions from exits during training. Our approach offers scientists a practical way to achieve both high performance and interpretability, addressing a fundamental challenge in machine learning for scientific discovery.
@@ 1 Introduction
Despite these promising qualities, KANs face challenges that can limit their effectiveness in scientific applications. Training KANs presents optimization difficulties, as the learnable univariate functions must be carefully parameterized and refined, with deeper networks often proving particularly challenging to optimize [15]. The architecture search problem—determining the appropriate number of layers and widths—also remains significant, as practitioners must balance expressiveness against parsimony while avoiding overfitting [20]. Seeking smaller, more parsimonious models without sacrificing accuracy is crucial because overly deep KANs lose interpretability as compositions of many univariate functions become difficult to understand, while smaller KANs better retain the interpretability that makes them valuable for scientific modeling [15]. These challenges suggest that architectural innovations are needed to better realize KANs’ potential for scientific discovery.
@@
Experiments demonstrate the effectiveness of multi-exit KANs across diverse scientific modeling tasks, showing consistent improvements in both accuracy and parsimony compared to traditional single-exit KANs. Our key contributions include: (1) the first application of multi-exit architectures to KANs, with a joint training framework that enables deep supervision; (2) empirical validation across regression problems, dynamical systems modeling, continual learning scenarios, and real-world datasets; (3) evidence that multi-exit KANs often achieve better performance at earlier exits, indicating more parsimonious models without sacrificing accuracy; and (4) a “learning-to-exit” algorithm that addresses the choice of extra hyperparameters when using multi-exits. Additionally, we provide insights into why multi-exit architectures are particularly well-suited for KANs and discuss the mechanisms underlying their improved performance. These results suggest that multi-exit architectures represent a valuable enhancement to KANs for scientific applications, offering a principled approach to balancing accuracy and interpretability.
The rest of this paper is organized as follows. Section 2 provides background on Kolmogorov–Arnold networks and multi-exit and early-exit neural architectures. Section 3 describes our approach for incorporating multiple exits into the KAN architecture. Section 4 presents results comparing single-exit and multi-exit KANs across multiple domains. Section 5 introduces the learning-to-exit method. Finally, we conclude in Sec 6 by discussing the implications of our findings, including why multi-exit architectures improve KAN performance, and directions for future work.
@@ 2.1 Kolmogorov–Arnold networks
In a KAN, the activation functions are typically parameterized using B-splines, local polynomial approximations of the one-dimensional functions. Although many other function-fitting techniques have been considered, including radial basis functions [30], Fourier series [31], sinusoidal functions [32], Chebyshev polynomials [33], and wavelet-based representations [34], all of which have many advantages, particularly in terms of computational efficiency, for our purposes here we focus on the original B-spline approach, though we also demonstrate the generalizability of our multi-exit architecture using Fourier series in App. B.
@@
To form a deep network by stacking layers of learned activation functions, summation units (and, optionally, multiplication [16]) aggregate the outputs from the previous layer’s activation functions. The number of units across $L$ layers is $d=N_{0},N_{1},\ldots,N_{L}=m]$. The shape of the KAN is the vector of widths $N_{0},N_{1},\ldots,N_{L}]$. (See Fig. 1 for an example of a KAN network with shape $[2,3,2,1]$.) Between layers $i$ and $i+1$ there are $N_{i}N_{i+1}$ activation functions. Following [15], $\phi_{\ell,j,i}(x)$ denotes the activation function connecting unit $i$ in layer $\ell$ to unit $j$ in layer $\ell+1$ ($\ell=0,\ldots,L-1$; $i=1,\ldots,N_{\ell}$; $j=1,\ldots,N_{\ell+1}$). The summation units then combine the activation functions to propagate information through the network: for the signal $x_{\ell+1,j}$ into unit $j$ in layer $\ell+1$, giving
@@
KANs are trained with a loss function $\mathcal{L}=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{reg}}$ comprising a data loss and a regularization loss, with the hyperparameter $\lambda$ determining regularization strength, with $\lambda=0$ corresponding to no explicit regularization. Usually the data loss is mean squared error (MSE):
@@
where $y_{i}$ is the true value for observation $i$ and $\hat{y}_{i}$ is the KAN’s prediction for that observation, while the regularization loss is a combination of an L1 norm and an entropy both defined on the activation functions (for details, see Liu et al. [15]). Activation function parameters are learned via gradient-based optimization to minimize $\mathcal{L}$ on training data. Training commonly includes a refinement process where the resolution of the B-spline grids is increased; gradually increasing grid resolution during training has been shown to improve KAN accuracy better than using a finer grid from the start [15]. This grid refinement process acts as a form of path regularization. In addition to refinement, Liu et al. [15] also introduce pruning of weak links as a post-training regularization step (we discuss the potential of pruning in Sec. 6). KANs are typically optimized with quasi-Newton methods, particularly L-BFGS [35], also employed in this work, though first-order methods such as SGD or Adam [36] can be used as well.
@@ 2.2 Multi-exit and early exit networks
Multi-exit networks are a deep learning architecture where some or all hidden layers in the network have an additional branch point called an exit [21, 22, 23]. These exits are small subnetworks, often only a single layer, whose output can be used alongside or instead of the main trunk network’s output. The most common application is for deep classifiers [21, 22, 37, 38], often for computer vision or language models. In a classification task, the network is typically given inputs of varying difficulties, For easy inputs, i.e., those far from the decision boundary, the network may be able to classify accurately using only basic features built by the earlier layers. But for difficult inputs, the network may need to utilize all the layers to build more complex features in order to make a successful classification [21]. By equipping the network with multiple exits, an easy input can reduce inference-time compute by using the output of an early exit only, saving both time and energy when making predictions [24]. The promise of efficiency gains motivates the study of early exit networks and learning-to-exit algorithms. This approach goes by various names in the literature, including deeply supervised networks, cascaded learning, conditional deep learning, and adaptive inference. For more on multi-exit and early-exit networks, see Scardapane et al. [24], Laskaridis et al. [39] or Rahmath et al. [40].
@@ 3 Adding exits to KANs
A multi-exit KAN augments a standard (single-exit) KAN of shape $d=N_{0},N_{1},\ldots,N_{L}=m]$ with additional exits as follows. For each layer $\ell$ of the KAN, beginning from the input and continuing until the second-to-last hidden layer [^1] The last hidden layer already has an exit, the main trunk output., add an exit layer, another KAN network, with shape $N_{\ell},m]$. The $m$-dimensional output of these exits can then be used to predict the same output as the main trunk exit. We illustrate a multi-exit KAN alongside the corresponding single-exit version in Fig. 1.
@@
where $\mathcal{L}_{k}=\sum_{i}\left(y_{i}-\hat{y}_{i}^{(k)}\right)^{2}/n$ is the MSE for exit $k$ and $w_{k}$ is the weight for exit $k$. The exit weights satisfy $\sum_{k}w_{k}=1$. These weights become a hyperparameter to be tuned by the researcher using validation data and this tuning process in practice has been straightforward. Equation (9) is quite flexible as the exit weights allow us to prioritize certain exits by weighting them more heavily than others, and individual exits can even be disabled by zeroing their weights. However, for a network with many exits, these K1K-1 degrees of freedom become a large search space. Therefore, after our main results, in Sec. 5 we propose and apply a “learning-to-exit” method to automatically learn $w$ alongside the KAN parameters.
@@
With more parameters requiring more training time, it is important to understand how many more parameters are added to a KAN by adding exits. The number of parameters in a KAN depends on its architecture, which dictates the number of activation functions, and the number of parameters per activation function. The parameters per activation function depends on the number and order of the spline bases (Eq. (5)), assuming B-splines are used to model the activation functions. This is the same for activation functions in the main trunk and in the exits, so we only need to consider the number of activation functions to determine the overhead added to a KAN by adding exits.
The number of activation functions between layers in a standard KAN is the product of the layer widths, so a KAN with shape $d=N_{0},N_{1},\ldots,N_{L}=m]$ will have
@@
Notice that the sum of layer widths will be smaller than the sum of products of adjacent layer widths, (unless the layers are all one unit), so, unless $m$ is large, the main trunk dominates the number of parameters in the KAN. Indeed, for the case of uniform layer width, $N_{\ell}=d$ for all $\ell$, the main trunk will have $(L-1)d^{2}+dm$ activation functions, including the original exit, and the newly added exits will introduce (L1)dm(L-1)dm activation functions in total. In this case, the new exits will contribute fewer activation functions than the main branch when $m<d$, typical of regression problems (Eq (1)) and a nearly equal number (fewer by $dm$) of activation functions when $m=d$, typical of dynamical systems modeling (Eqs. (2) or (3)).
Also, note that no one prediction made by the model will use all the activation functions, even if they were all used during training. Thus, multi-exits provide their benefits with reasonable, often modest, parameter overhead. Training time overhead is similarly modest, as detailed in Appendix A.
@@ 4 Results
Experiments compare single-exit and multi-exit KANs on various regression tasks of known functional forms (Sec. 4.1; Figs. 2 and 3; Table 1), on multi-step forecasting of dynamical systems (Sec. 4.2; Figs. 4 and 5), on a model of continual learning (Sec. 4.3, Fig. 6), and on three real-world datasets ((Sec. 4.4, Table 2). For each task, a manual architecture search identifies good KAN shapes and exit weights, while holding all other hyperparameters fixed. Performance is assessed with the root mean squared error ($\mathit{RMSE}$) of its predictions on test data (as well as $R^{2}$ values for the real-world data). Regarding the exit weights, Sec. 5 explores learning the weights automatically with a “learning-to-exit” approach.
@@ 4.1 Regression tasks
For the sinc function a KAN of shape $[1,2,2,2,1]$ performed well. See Appendix A for full details on training settings and hyperparameters and data generation. This KAN may seem deep for such a function. KANs can support multiplication units [16], although they are not strictly necessary due to the KART, so we decided to forgo them for simplicity and therefore expected a deeper KAN to perform better for this task, hence the aforementioned shape. Parameterizing the multi-exit weighted loss with a simple linear ramp, $w=[1,2,3,4]$ (unnormalized) worked well: performance on test data was good, with the final exit (shown in Fig. 2) showing an order of magnitude lower error than the equivalent single-exit network. In fact, three of the four exits, all but Exit 0, outperformed the single-exit network, indicating robust and parsimonious (parameter-efficient) capture of the data.
(Note that neither model shown in Fig. 2 is optimal, a point we return to in Sec. 6 as it illustrates important aspects of using a multi-exit architecture.)
@@
In the multi-exit KAN, unsurprisingly, the initial exit, which lacks any composability, is unable to represent this function. The next exit, however, does well, achieving error one fourth that of the larger, single-exit model ($\mathit{RMSE}=0.0045$ vs. $0.0232$). The final exit at $\mathit{RMSE}=0.007$ also outperforms the same-size single-exit model. This result demonstrates that multi-exit KANs can identify the right level of complexity, with the middle exit outperforming both simpler and more complex alternatives.
@@
Each equation was converted to the dimensionless form indicated in the table and used to generate data (see Appendix A). KANs with shape $[d,5,5,5,5,1]$) were fitted to each dataset. The multi-exit KAN found good results with $w=[0,0,1,1,3/2]$ (only Exits 2–4 are active) on Eq. I.27.6, and we proceeded to use this $w$ across all equations. For each multi-exit KAN, Table 1 reports the smallest $\mathit{RMSE}$ across its exits.
Feynman Eq. Original Formula Dimensionless formula # vars $\mathit{RMSE}$ (single) $\mathit{RMSE}$ (multi) (Exit)
@@
As shown in Table 1, multi-exit networks achieved lower test $\mathit{RMSE}$ than the single-exit model in nine of ten cases. Interestingly, and in line with our observations from the 2D regression shown in Fig. 3, for half of the problems, the best performing exit is not the final exit, indicating we find more parsimonious (smaller) models with multi-exits that are also more accurate than larger, single-exit models.
@@ 4.2 Data-driven modeling of dynamical systems
where $\phi_{n}=0.4-6\left(1+x_{n}^{2}+y_{n}^{2}\right)^{-1}$ and bifurcation parameter $\mu=0.9$. KANs, unlike sparse regression methods, have been shown to model the Ikeda map well [19].
KANs found good results modeling the Ikeda map with a $[2,4,4,4,2]$ shape and, for the multi-exit KAN, exit weights $w=[0,0,1,2]$ (the first two exits were disabled). (Note that Panahi et al. [19] used a fixed grid $G=10$ while we used grid refinement (see Appendix) for both single- and multi-exit KANs.) For one-step prediction the multi-exit KAN achieved $\mathit{RMSE}=4.560\times 10^{-3}$ compared to the single-exit’s $5.484\times 10^{-3}$, an improvement of $16.8\%$. For multi-step prediction, as shown in Fig. 4 the multi-exit KAN tracks the dynamics for approximately twice as many timesteps as the single-exit KAN before inevitably diverging due to chaos.
Figure 4: Multi-step prediction of the Ikeda map (Eq. (14)). The multi-exit KAN tracks the dynamics well for about twice as many steps into the future (shaded regions) as the single-exit KAN. (Blue: ground truth; orange: KAN prediction.)
@@
KANs performed well modeling the ecosystem with shape $[3,3,3,3]$ KANs and $w=[2,1,1/2]$ for the multi-exit KAN, achieving very good multi-step prediction (Fig. 5) but slightly worse one-step prediction ($\mathit{RMSE}=3.774\times 10^{-4}$ for the multi-exit compared to $3.171\times 10^{-4}$ for the single-exit KAN). The single-exit KAN, while still performing well, appears to be overfit to the local trajectory while the multi-exit KAN better captured the underlying attractor structure.
Figure 5: Multi-step prediction of the ecosystem (Eq. (15)). The multi-exit KAN encodes the dynamics well, not diverging significantly until $t>1000$.